OpenStudy (anonymous):

If log10(x-2) - 3log10(2x)=1-log10(y) then y is equal to? Working out please :D The answer is (80x^3)/x-2

OpenStudy (lalaly):

do u know the rules of logarithms?

OpenStudy (lalaly):

when u have log a - log b = log (a/b) and alogb = logb^a so here log10 (x-2) - log10 (2x)^3 = 1-log10 y it becomes log10 (x-2)/(2x)^3 = 1-log10 y

OpenStudy (lalaly):

and we also know that log a + log b = log a*b so log10 (x-2)/(2x)^3 +log10 y =1 so log10 ((x-2)/8x^3)y =1 10^1 =( x-2 / 8x^3)y y= 80x^3 / x-2

OpenStudy (anonymous):

Thanks for your help. Yeah i need to brush up on my log rules xD

OpenStudy (lalaly):

yur welcome :D